The electric motor of a model train accelerates th
Answer #1

here,

final speed , v = 0.68 m/s

time taken , t = 0.031 s

let the accelration be a

v = u +a * t

0.68 = 0 + a * 0.031

a = 21.94 m/s^2

force , F = m * a

F = 0.555 * 21.94 N

F = 12.17 N

average velocity , v = (0.68 - 0)/2 = 0.34 m/s

the average power , P = F * v

P = 12.17 * 0.34

P = 4.14 W

the average power delivered is 4.14 W